OpenStudy (anonymous):

MEDALS (geometry) Determine whether R(3, 2), O(6, 3), and Q(6, 2) can be the vertices of a triangle.

OpenStudy (solomonzelman):

find the distances of the sides

OpenStudy (anonymous):

the question what the shortest side to o to line OR if OQ perpendicular to QR

OpenStudy (solomonzelman):

\[d=\sqrt{(y_1-y_2)^2+(x_1-x_2)^2}\] distance formula. plug in each two points to find the distance btw them.

OpenStudy (solomonzelman):

Woops, misread it.

OpenStudy (anonymous):

what ? so do i not do that ?

OpenStudy (anonymous):

?

OpenStudy (solomonzelman):

No, I think I read it write. So you do do the distance formula. Draw the triangle with the sides and label the sides. (For your first step)

OpenStudy (anonymous):

a. no: QR + RO = QO distance ≈ 1 b. no: QR + QO = RO distance ≈ 4 c. yes: QR + RO < QO QR + QO < RO QO + RO < QR distance ≈ 4 d. yes: QR + RO > QO QR + QO > RO QO + RO > QR distance ≈ 1 these are the answers

OpenStudy (anonymous):

just to show you what i need to find

OpenStudy (solomonzelman):

Don't do ≈, give me the exact value, like sqrt of. also did you mean vertex (instead of vertices)

OpenStudy (anonymous):

im confused, i dont know how to wok out the problem i know how to do the distance thing but that dosent give me my answer

OpenStudy (solomonzelman):

Yeah, and I am kind of a bad helper too (and bad mathmatician). But, anyway... We have a triangle. O-R=(3,1) and Q-R=(3,0). These are linearly independent. So clearly it is a triangle, not just a line. OQ=Q-O=(0,-1) and from before QR=-RQ=(-3,0). We see that these sides are perpendicular as 'OQ,QR' =-3*0-1*0=0 and because OQ is perpendicular to QR, it is the shortest line from O to QR. Therefore the shortest distance is the length of OQ which is 1.